OpenStudy (anonymous):

The half- life of- Barium- 137m is 2.552 minutes. how much of an 18 gram sample would be left after 10 minutes

OpenStudy (wolf1728):

Ending Amount = Beginning Amount / 2^n where 'n' is elapsed time / half-life n = 10 / 2.552 n = 3.918495297806 Ending Amount = 18 / 3.918495297806 INCORRECT Ending Amount = 4.5936 INCORRECT Calculator to check your work and formulas: http://www.1728.org/halflife.htm

OpenStudy (wolf1728):

Ending Amount = 18 / 2^3.918495297806 Ending Amount = 18 /15.1211430692 Ending Amount = 1.1903861975 Calculator to check your work (which I should have done) http://www.1728.org/halflife.htm